ChemCo is a price-taker able to receive $10 for each bottle of vitamins they sell. At the profit maximizing level of production their AVC = $5 and AFC = $3. Given the following points on their MC curve, what is their maximum possible profit?

Quantity 0 100 200 300 400 500 600
Marginal Cost n/a $8 $4 $6 $8 $10 $15

        

Group of answer choices

No profit

$4000

$5000

$1000

$500
Answer #1

Profit maximizing level of output determined by P=MC

For P=10=MC firm will produce Q=500 units

Total revenue =P*Q=10*500=5000

Total cost =(AVC+AFC)*Q=(5+3)*500=4000

Profit =Total revenue-Total cost=5000-4000=$1000

Answer: $1000